OpenStudy (usukidoll):

First order ode... dr/dt=a/p-br/3p

OpenStudy (usukidoll):

\(\frac{dr}{dt}=\frac{a}{p}-\frac{b}{3p}r\)\\ Standard Linear ODE form:\\ \(\frac{dy}{dx}+p(x)y=q(x)\)\\ Since we have \(y=r\) and \(x=t\), we have\\ \(\frac{dr}{dt}+p(t)r=q(t)\)\\ Therefore, our \(p(t) = \frac{b}{3p}\) and our \(q(t) = \frac{a}{p}\)\\ Also, our \(v(t) = e^{\int \frac{b}{3p}} \rightarrow e^{ \frac{b}{3p}t}\)\\ Multiplying our integrating factor throughout the equation, we have, \\ \(e^{\frac{b}{3p}t} (\frac{dr}{dt}) + e^{ \frac{b}{3p}t} (\frac{b}{3p}) =e^{\frac{b}{3p}t}(\frac{a}{p})\)\\ Using reverse product rule, we have,\\ \(r (e^{ \frac{b}{3p}t}) =\int e^{ \frac{b}{3p}t}(\frac{a}{p})\)\\ Integrating the right side, we have\\ \(r (e^{ \frac{b}{3p}t}) =\frac{3a}{b}(e^{ \frac{b}{3p}t})+C \)\\ Dividing \((e^{ \frac{b}{3p}t})\) throughout the whole equation, we have,\\ \(r =\frac{3a}{b}+\frac{C}{(e^{ \frac{b}{3p}t})} \)\\ \(r =\frac{3a}{b}+C((e^{- \frac{b}{3p}t})) \)\\ Now, suppose our Initial Condition is \(r(0)=0\). Then we have,\\ \(r(0) =\frac{3a}{b}+C((e^{- \frac{b}{3p}(0)})) \)\\ \(0 =\frac{3a}{b}+C(e^{0}) \)\\ \(0 =\frac{3a}{b}+C \)\\ \(-\frac{3a}{b}=C \)\\ Substituting C back into the equation, we have, \\ \(r =\frac{3a}{b}-\frac{3a}{b}((e^{- \frac{b}{3p}t})) \)\\

OpenStudy (usukidoll):

I think the fact that using separation of variables is impossible on this due to the fact that it's not in the form of h(y) dy =f(x) dx . Our equation is \(\frac{dr}{dt}=\frac{a}{p}-\frac{b}{3p}r\)\\. We can only use separation of variables with multiplication or division, but never with addition or subtraction.

OpenStudy (usukidoll):

like if our equation was \[/\frac{dx}{dy} = x\] we can use separation of variables, but the equation I was given breaks the rules real fast.

OpenStudy (usukidoll):

you can't have something like \[\frac{dy}{dx} = x+y \]

OpenStudy (usukidoll):

and expect separation of variables to work.. you can't add to both sides or subtract both sides XD

ganeshie8 (ganeshie8):

dr/dt=a/p-br/3p is separable right ?

OpenStudy (usukidoll):

I don't think so? Not unless the r is everywhere on the right side

OpenStudy (usukidoll):

I have this equation  there is no () around